Identify the values of coefficients a,b, and c in each equation. 1. 5x2-6x+11=0 a=__b=__c=__ 2. x2=3x+7 a=__b=__c=__
Answer by Edwin McCravy(20065)

Identify the values of coefficients a,b, and c in each equation.

1. 5x²-6x+11=0   a=__b=__c=__
    5  -6  11
    a   b   c     a=5, b=-6, c=11

2.      x²=3x+7  <-- must get 0 on the right side, left side in descending order.
   x²-3x-7=0     <-- put a 1 before the x²
  1x²-3x-7=0  
  1  -3 -7
  a   b  c       a=1, b=-3, c=-7 


Edwin